# SingingHead: A Large-scale 4D Dataset for Singing Head Animation

### v2

@@ 3.2 Data Processing
Figure 3: Overview of the UniSinger framework. Given singing audio and shape parameter, the 3D facial animation module can generate diverse singing facial motions synchronized with the input singing signal. The generated 3D facial motion can be further rendered to 2D singing video through the face renderer module.
4 Method
@@
Figure 4: VAE-based 3D facial animation. Given the 3D facial motion sequence $\{\boldsymbol{m}_{1},...,\boldsymbol{m}_{T}\}$, FLAME shape parameter $\boldsymbol{\beta}$ and singing audio $\boldsymbol{A}_{1:T}$, the transformer-based VAE encoder outputs the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$. By sampling from the distribution, latent vector $\boldsymbol{z}$ is obtained and fed into the transformer-based VAE decoder together with the shape parameter and input audio, and output the generated facial motion sequence $\{\boldsymbol{\hat{m}}_{1},...,\boldsymbol{\hat{m}}_{T}\}$. During inference time, we can generate diverse facial motion sequences according to the input audio and shape parameter by sampling the learned distribution via the VAE decoder.
@@
Along with the SiningHead dataset, we argue that 3D and 2D facial animation tasks can be solved together in one pipeline. To this end, we propose a unified framework called UniSinger to generate diverse 3D facial motion according to the input singing audio (Sec.4.1), which can be further rendered into 2D portrait singing videos (Sec.4.2). At inference time, given a single image and singing audio, we first reconstruct the FLAME [27] shape parameters $\boldsymbol{\beta}$ and camera parameter $\boldsymbol{c}$ from the input image through a monocular 3D reconstruction method [9]. Then, we feed the audio and shape parameter into the 3D facial animation module to get the 3D singing motion represented by a sequence of FLAME expression $\boldsymbol{\psi}$ and pose $\boldsymbol{\theta}$ parameters. The generated 3D facial can be directly used for virtual avatar animation, or further rendered to depth video and fed into our face renderer together with the single input image to get a 2D portrait singing video. The overview of our pipeline is illustrated in Fig. 3.
4.1 VAE-based 3D Facial Animation
@@
Given audio $\boldsymbol{A}_{1:T}=\{\boldsymbol{a}_{1},...,\boldsymbol{a}_{T}\}$ and FALME shape parameter $\boldsymbol{\beta}\in\mathbb{R}^{100}$, our VAE-based 3D facial animation module can generate diverse 3D facial motion $\boldsymbol{M}_{1:T}=\{\boldsymbol{m}_{1},...,\boldsymbol{m}_{T}\}$, where $\boldsymbol{m}_{i}=(\boldsymbol{\psi}_{i},\boldsymbol{\theta}_{i})$ denotes the motion of $i$-th frame which is represented by FLAME expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$ and pose $\boldsymbol{\theta}\in\mathbb{R}^{50}$ parameters. Concretely, we utilize a conditional variational autoencoder (CVAE) model [45] inspired by [39] to construct our 3D facial animation module, which is built on transformer layers [49] and conditioned on shape and audio embeddings (Fig. 4).
@@
Transformer-based VAE encoder. The VAE encoder takes a facial motion sequence and the corresponding shape parameter and singing audio as input, and outputs the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ of the singing facial motion space. Then a ${d}$-dimensional latent vector $\boldsymbol{z}\in\mathbb{R}^{d}$ is sampled from the Gaussian distribution using reparameterization [25], which will be fed to the VAE decoder. To be specific, the facial motion sequence is first projected into a ${d}$-dimensional space via a linear projection layer. To model the style of facial motion of different people during singing, the FLAME shape parameter, which varies from person to person, is embedded to a ${d}$-dimensional feature vector via the shape embedding layer. As for the input audio, we build an audio encoder using the structure of a SOTA speech model, wav2vec 2.0 [2] followed by a linear projection layer to project the output audio feature to the $d$ dimension following [12, 53], and initialize our audio encoder using the pre-trained wav2vec 2.0 weights. Since the number of tokens of remains unchanged across multiple transformer layers, we add two additional learnable tokens $[\mu]$ and $[\sigma]$ to extract the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ inspired by [10, 39]. After these steps, all feature dimensions are unified to a ${d}$-dimensional format. We add shape embedding to the projected motion sequence, and then feed them together with the learnable tokens into a transformer decoder [49] after applying a periodic positional encoding (PPE) [12] to all the input tokens. At the same time, the audio feature is feed into the transformer decoder through cross-attention layers with a diagonal attention mask to align the audio with facial motion sequence. Please refer to the Sup. Mat. for more details. Finally, the first two tokens of the transformer outputs are picked out as the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$.
@@
Transformer-based VAE decoder. Given $T$ latent vectors $\{\boldsymbol{z},...,\boldsymbol{z}\}$, together with shape parameter and audio, the VAE decoder generates a $T$-frames 3D facial motion sequence according to the input audio. Concretely, the shape parameter and audio are embedded to the same way as in the VAE encoder, and PPE [12] is applied to the shape embedding and $T$ latent vectors. Then, the output of PPE is fed into another transformer decoder together with the audio feature through cross-attention layers similar to the VAE encoder. The transformer outputs without the first toke is finally fed to a linear projection layer to get the generated 3D facial motion $\boldsymbol{\hat{M}}_{1:T}=\{\boldsymbol{\hat{m}}_{1},...,\boldsymbol{\hat{m}}_{T}\}$.
@@
Training and inference. We train the VAE-based 3D facial animation module using the audio, shape parameter, and facial motion pairs in our SingingHead Dataset. We use the loss function $\mathcal{L}=\lambda_{r}\mathcal{L}_{rc}+\lambda_{v}\mathcal{L}_{vel}+\lambda_{k}\mathcal{L}_{kl}$ to train our model, where $\mathcal{L}_{re}$, $\mathcal{L}_{vel}$, $\mathcal{L}_{kl}$ stands for reconstruction loss, velocity loss and KL loss, respectively. Specifically, reconstruction loss measures the distance between the generated facial motion $\boldsymbol{\hat{M}}_{1:T}$ and the ground truth facial motion $\boldsymbol{M}_{1:T}$, which is written as:
@@
KL loss $\mathcal{L}_{kl}$ aims at regularizing the latent distribution to a standard Gaussian distribution.
@@
During inference time, only the VAE decoder is used. Concretely, given audio, FLAME shape parameter, and the latent vector $\boldsymbol{z}$ randomly sampled from Gaussian distribution, our VAE decoder can generate realistic and accurate 3D singing facial motion synchronized with the input audio. Moreover, our model can generate diverse facial motions with the different sampling results of $\boldsymbol{z}$.
4.2 Face Renderer
@@
Take the FLAME depth image sequence and a reference image as input, our face renderer can render it to a portrait singing video with the same identity as the reference image. Specifically, the face renderer is actually a generator constructed based on a U-Net structure with skip connections following [33, 30]. Inspired by [30], temporally positional embedding (TPE) is concatenated with the input depth image sequence and reference image and fed into the generator to improve temporal stabilization. The generator is trained together with a PatchGAN [19] based discriminator through generative adversarial training strategy and multiple losses including GAN loss, photo loss, perceptual loss, and feature matching loss similar to [30].
5 Experiments
Figure 5: Qualitative comparison with 3D facial animation methods. We show the generated facial motion sequences according to two input audios. UniSinger can generate more accurate and natural sequences compared with the other two methods.
5.1 Experimental Setup
@@
Implementation details. We train the proposed UniSinger framework on our SiningHead dataset. Specifically, we cut the sequences in the dataset into equal-length 8s segments, and delete the cropped segments less than 8s in length, which leads to a set of data containing 12196 sequences. Then, the dataset is split into training, validation, and test sets with a ratio of $80\%:5\%:15\%$. We use all the 9758 sequences in the training set to train the VAE-based 3D facial animation module, and test it on the whole test set which contains 1829 sequences. The training process takes about 9 hours on 4 NVIDIA 3090 GPUs. As for the face renderer, we train a separate network for each person similar to [30, 56, 16]. We select 10 identities for evaluation. For each identity, the training of the face renderer takes about 10 hours on a single NVIDIA 3090 GPU.
Evaluation metrics. We evaluate the 3D singing head animation and 2D singing portrait animation respectively using the following metrics. For 3D singing head animation, we define three metrics to measure the accuracy and variety of the generated 3D facial motion:
  • Minimum distance (MinDist), i.e., the minimum value of the distances between multiple generated facial motions and the ground truth, which is computed as $\underset{\boldsymbol{M}\in X}{min}||\boldsymbol{\hat{M}}-\boldsymbol{M}||_{2}^{2}$, where $X=\{\boldsymbol{M}_{1},...,\boldsymbol{M}_{n}\}$ includes all the generated faces and $n$ is the sampling number.
  • Mean distance (MeanDist), i.e., the average of the distances between multiple generated facial motions and the ground truth, computed as $\frac{1}{n}{\textstyle\sum_{i=1}^{n}}||\boldsymbol{\hat{M}}_{i}-\boldsymbol{M}_{i}||_{2}^{2}$.
  • Average Pairwise Distance (APD). Inspired by yuan2020dlow, we use the average L2 distance between all pairs from the generated $n$ samples of facial motions to measure the diversity within samples, which is computed as $\frac{1}{n(n-1)}{\textstyle\sum_{i=1}^{n}}{\textstyle\sum_{j\neq i}^{n}}||\boldsymbol{M}_{i}-\boldsymbol{M}_{j}||_{2}^{2}$.
For 2D singing portrait animation, we evaluate the quality of the generated video and whether the facial movements are synchronized with the input audio. We use Frechet Inception Distance (FID) [37, 43] and SSIM [52] to evaluate the quality of the generated video frames. As for the correctness of the generated video, we use landmark distance (LMD) [6] of the mouth region between the generated and the ground truth video frames to evaluate whether the mouth is animated correctly according to the input audio.
5.2 Compare with 3D Facial Animation Methods
Our Unisinger framework can achieve audio-driven 3D facial animation through the VAE-based 3D facial animation module as shown in Fig. 4. To demonstrate the superiority of the 3D facial animation module of UniSinger, we compare it with two SOTA 3D facial animation methods: FaceFormer [12] and CodeTalker [53]. Note that all the methods are trained on the SingingHead training set and tested on the whole SingingHead testing set using its 3D facial motion data. The qualitative results and qualitative results are shown in Fig. 5 and Tab. 2 respectively. We can see that our method achieves significantly lower MinDist and MeanDist, which shows its powerful ability to generate accurate and synchronized 3D facial motions according to the input audio, which can also be seen from Fig. 5. Specifically, our method can generate more accurate facial movements especially in the mouth region, while other methods can hardly model the facial motions and tend to produce static results or meaningless shaking.
Method MinDist $\downarrow$ MeanDist $\downarrow$ APD $\uparrow$
FaceFormer [12] 4.064 4.064 -
CodeTalker [53] 6.876 6.876 -
Ours (n=1) 2.859 2.859 -
Ours (n=10) 2.677 2.862 10.366
Ours (n=30) 2.640 2.860 10.369
Table 2: Comparison with 3D facial animation methods. We report the minimum distance, mean distance, and average pairwise distance of the generated 3D facial motion on the SingingHead testing set. Note that only the generation results with diverse samples have the APD value.

### v3

@@ IV Method
Figure 6: Overview of the UniSinger framework. Given singing audio, background music, and shape parameter, the 3D facial animation module can generate diverse singing facial motions synchronized with the input singing signal. The generated 3D facial motions can be further rendered into 2D singing videos through the face renderer module.
@@
Figure 7: VAE-based 3D facial animation. Given the 3D facial motion sequence $\{\boldsymbol{m}_{1},...,\boldsymbol{m}_{T}\}$, FLAME shape parameter $\boldsymbol{\beta}$, singing audio $\boldsymbol{A}_{1:T}$ and background music $\boldsymbol{B}_{1:T}$, the transformer-based VAE encoder outputs the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$. By sampling from the distribution, latent vector $\boldsymbol{z}$ is obtained and fed into the transformer-based VAE decoder together with the shape parameter and input audio, and output the generated facial motion sequence $\{\boldsymbol{\hat{m}}_{1},...,\boldsymbol{\hat{m}}_{T}\}$. During inference time, we can generate diverse facial motion sequences according to the input audio and shape parameter by sampling the learned distribution via the VAE decoder.
@@
Along with the SingingHead dataset, we argue that 3D and 2D facial animation tasks can be solved together in one pipeline. To this end, we propose a unified framework called UniSinger to generate diverse 3D facial motion according to the input singing audio (Sec.IV-A), which can be further rendered into 2D portrait singing videos (Sec.IV-B). At inference time, given a single image, singing audio, and background music, we first reconstruct the FLAME [35] shape parameters $\boldsymbol{\beta}$ and camera parameter $\boldsymbol{c}$ from the input image through a monocular 3D reconstruction method [77]. Then, we feed the audio and shape parameter into the 3D facial animation module to get the 3D singing motion represented by a sequence of FLAME expression $\boldsymbol{\psi}$ and pose $\boldsymbol{\theta}$ parameters. The generated 3D facial can be directly used for virtual avatar animation, or further rendered to depth video and fed into our face renderer together with the single input image to get a 2D portrait singing video. The overview of our pipeline is illustrated in Fig. 6.
IV-A VAE-based 3D Facial Animation
@@
Given singing audio $\boldsymbol{A}_{1:T}=\{\boldsymbol{a}_{1},...,\boldsymbol{a}_{T}\}$, background music $\boldsymbol{B}_{1:T}=\{\boldsymbol{b}_{1},...,\boldsymbol{b}_{T}\}$ and FLAME shape parameter $\boldsymbol{\beta}\in\mathbb{R}^{100}$, our VAE-based 3D facial animation module can generate diverse 3D facial motion $\boldsymbol{M}_{1:T}=\{\boldsymbol{m}_{1},...,\boldsymbol{m}_{T}\}$, where $\boldsymbol{m}_{i}=(\boldsymbol{\psi}_{i},\boldsymbol{\theta}_{i})$ denotes the motion of $i$-th frame which is represented by FLAME expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$ and pose $\boldsymbol{\theta}\in\mathbb{R}^{50}$ parameters. Concretely, we utilize a conditional variational autoencoder (CVAE) model [81] inspired by [38] to construct our 3D facial animation module, which is built on transformer layers [36] and conditioned on shape and audio embeddings (Fig. 7).
@@
Transformer-based VAE encoder. The VAE encoder takes a facial motion sequence and the corresponding shape parameter and singing audio as input, and outputs the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ of the singing facial motion space. Then a ${d}$-dimensional latent vector $\boldsymbol{z}\in\mathbb{R}^{d}$ is sampled from the Gaussian distribution using reparameterization [37], which will be fed to the VAE decoder. To be specific, the facial motion sequence is first projected into a ${d}$-dimensional space via a linear projection layer. To model the style of facial motion of different people during singing, the FLAME shape parameter, which varies from person to person, is embedded to a ${d}$-dimensional feature vector via the shape embedding layer. As for the input singing audio, we build an audio encoder using the structure of a SOTA speech model, wav2vec 2.0 [82] followed by a linear projection layer to project the output audio feature to the $d$ dimension following [4, 5], and initialize our audio encoder using the pre-trained wav2vec 2.0 weights. Considering that background music mainly provides rhythmic and beat information instead of phonemes, we utilize its low-level MFCC features [83, 84]. Then, a linear projection layer is used to project the 64-dimensional MFCC features to the $d/2$ dimension, which is then concatenated with the singing audio feature to obtain the entire audio feature. Finally, the entire audio feature is projected to $d$ dimension using another linear projection layer. Since the number of tokens remains unchanged across multiple transformer layers, we add two additional learnable tokens $[\mu]$ and $[\sigma]$ to extract the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ inspired by [85, 38]. After these steps, all feature dimensions are unified to a ${d}$-dimensional format. We add shape embedding to the projected motion sequence, and then feed them together with the learnable tokens into a transformer decoder [36] after applying a periodic positional encoding (PPE) [4] to all the input tokens. At the same time, the entire audio feature is fed into the transformer decoder through cross-attention layers with a diagonal attention mask to align the audio with the facial motion sequence. Finally, the first two tokens of the transformer outputs are picked out as the distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$.
@@
Transformer-based VAE decoder. Given $T$ latent vectors $\{\boldsymbol{z},...,\boldsymbol{z}\}$, together with shape parameter and audio, the VAE decoder generates a $T$-frames 3D facial motion sequence according to the input audio. Concretely, the shape parameter and audio are embedded in the same way as in the VAE encoder, and PPE [4] is applied to the shape embedding and $T$ latent vectors. Then, the output of PPE is fed into another transformer decoder together with the entire audio feature through cross-attention layers similar to the VAE encoder. The transformer outputs without the first toke is finally fed to a linear projection layer to get the generated 3D facial motion $\boldsymbol{\hat{M}}_{1:T}=\{\boldsymbol{\hat{m}}_{1},...,\boldsymbol{\hat{m}}_{T}\}$.
@@
Training and inference. We train the VAE-based 3D facial animation module using the audio, shape parameter, and facial motion pairs in our SingingHead Dataset. We use the loss function $\mathcal{L}=\lambda_{k}\mathcal{L}_{kl}+\lambda_{r}\mathcal{L}_{rc}+\lambda_{v}\mathcal{L}_{vel}+\lambda_{m}\mathcal{L}_{mesh}$ to train our model, where $\mathcal{L}_{vel}$, $\mathcal{L}_{rc}$, $\mathcal{L}_{kl}$, $\mathcal{L}_{mesh}$ stands for KL loss, reconstruction loss, velocity loss and mesh vertices loss, respectively. Specifically, KL loss $\mathcal{L}_{kl}$ aims at regularizing the latent distribution to a standard Gaussian distribution, which is the same as the standard VAE [37]. Reconstruction loss measures the distance between the generated facial motion $\boldsymbol{\hat{M}}_{1:T}$ and the ground truth facial motion $\boldsymbol{M}_{1:T}$, which is written as:
@@
Mesh vertices loss measures the L2 distance between the vertices of the face region from the generated and ground truth mesh sequences. The meshes are obtained by applying the FLAME model [35] to the generated or ground truth FLAME parameters.
@@
During inference time, only the VAE decoder is used. Concretely, given input audio, FLAME shape parameter, and the latent vector $\boldsymbol{z}$ randomly sampled from Gaussian distribution, our VAE decoder can generate realistic and accurate 3D singing facial motion synchronized with the input audio. Moreover, our model can generate diverse facial motions with the different sampling results of $\boldsymbol{z}$.
Figure 8: Structure of the Face Renderer. Given a depth video that represents a singing motion and a reference image, the face renderer module can generate a singing portrait video through a UNet-like convolutional neural network with residual blocks.
IV-B Face Renderer
@@
As shown in Fig. 8, the face renderer module takes the FLAME depth image sequence and a reference image as input, rendering it to a portrait singing video with the same identity as the reference image. Specifically, the face renderer module is actually a generator constructed based on a convolutional neural network with skip connections following [86, 2]. In detail, the encoder of the generator contains 8 layers with resolutions of $(256^{2},128^{2},64^{2},32^{2},16^{2},8^{2},4^{2},2^{2})$ and corresponding feature channel numbers of $(64,128,256,512,512,512,512,512)$. The structure of the decoder is symmetric to the encoder, with upsampling layers. Inspired by [2], temporally positional embedding (TPE) is concatenated with the input depth image and reference image and then fed into the generator to improve temporal stabilization. The generator is trained together with a PatchGAN [87] based discriminator through generative adversarial training strategy and multiple losses including GAN loss, photo loss, perceptual loss, and feature matching loss similar to [2].
V Experiments
TABLE II: Quantitative results of 3D singing head animation. We report the evaluation metrics of the pre-trained models and the models trained on the SingingHead dataset. The units of LVE, FVE, and FDD are $mm$, $mm$, and ×$\times 10^{-2}mm$, respectively. The metrics of FaceDiffuser are calculated under 10 samples. For the 3D part of the proposed UniSinger framework (Ours), we report the results under 1, 10, and 30 samples, respectively. Note that only methods that generate head pose have the BA and PPE metrics, and only non-deterministic generation methods have the Diversity metric.
Method LVE-Mean/Min $\downarrow$ FVE-Mean/Min $\downarrow$ FDD $\downarrow$ BA $\uparrow$ PPE $\downarrow$ Diversity $\uparrow$ FPS $\uparrow$
FaceFormer [4] 13.025 / 13.025 2.8999 / 2.8999 2.2496 - - - 1.6171
CodeTalker [5] 13.103 / 13.103 2.9212 / 2.9212 2.1548 - - - 0.2209
Imitator [12] 12.591 / 12.591 2.8483 / 2.8483 2.9012 - 1.8329
Pre-trained SelfTalk [13] 12.971 / 12.971 2.9018 / 2.9018 1.5410 - - - 28.103
FaceFormer [4] 4.4715 / 4.4715 1.0255 / 1.0255 2.6241 - - - 1.4500
CodeTalker [5] 5.9122 / 5.9122 1.3406 / 1.3406 2.4946 - - - 0.2204
Imitator [12] 4.4188 / 4.4188 1.0265 / 1.0265 2.7503 - - - 1.9076
SelfTalk [13] 4.7316 / 4.7316 1.0963 / 1.0963 1.2911 - - - 27.360
FaceDiffuser [14] 8.8594 / 7.2877 2.1219 / 1.0907 1.1435 0.2152 0.2216 3.9516 0.0615
Ours (n=1) 4.5802 / 4.5802 1.0412 / 1.0412 1.2235 0.2382 0.0984 - 50.849
Ours (n=10) 4.5928 / 4.2806 1.0435 / 0.7909 1.2286 0.2378 0.0988 1.0993 50.849
Trained on SingingHead Dataset Ours (n=30) 4.5958 / 4.2242 1.0440 / 0.7232 1.2288 0.2381 0.0989 1.1006 50.849
V-A Implementation details
@@
We train the proposed UniSinger framework on our SiningHead dataset. Specifically, we cut the sequences in the dataset into equal-length 8s segments, and delete the cropped segments less than 8s in length, which leads to a set of data containing 12196 sequences. Then, the dataset is split into training, validation, and test sets with a ratio of $80\%:5\%:15\%$. We use all the 9758 sequences in the training set to train the VAE-based 3D facial animation module, and test it on the whole test set which contains 1829 sequences. The training process takes about 2 days on 4 NVIDIA 3090 GPUs. As for the face renderer, we train a separate network for each person similar to [2, 27, 64]. We select 10 identities for evaluation. For each identity, the training of the face renderer takes about 10 hours on a single NVIDIA 3090 GPU.
Moreover, we benchmark existing audio-driven 3D facial animation methods and 2D talking head methods on the SingingHead dataset. For the 3D singing head animation, we benchmark the pre-trained models [4, 5, 12, 13] using all the 1829 sequences in the SingingHead test set. We also train the existing methods [4, 5, 12, 13, 14] on the SingingHead training set, and report the results. For 2D singing portrait video synthesis, we use a total of 267 video sequences from the selected 10 identities as the test set, and benchmark some typical pre-trained 2D talking head models. Since our UniSinger framework can handle both 2D and 3D singing head animation tasks, we report its results on 3D singing head animation and 2D singing portrait video synthesis, respectively.
V-B 3D Singing Head Animation
Evaluation metrics. We evaluate the 3D singing head animation performance from four aspects. (1) Facial accuracy. Following existing audio-driven 3D facial animation methods, we measure the difference between the generated and the ground truth mesh vertices. To measure the lip synchronization, we adopt Lip Vertex Error (LVE) [18], which is calculated as the average value of the maximal $L_{2}$ error of all lip vertices in each frame. To evaluate the realism of the upper-face dynamics, we calculate the Upper-face Dynamics Deviation (FDD) similar to CodeTalker [5]. Moreover, we calculate the average $L_{2}$ distance of the generated and ground truth mesh vertices in the face region, termed Face Vertex Error (FVE), to measure the accuracy of facial area movements. Considering the non-deterministic methods, we report both the mean and minimum values of the LVE and FVE over all the generated samples. Note that all these metrics are calculated on the mesh vertices without head pose. (2) Head pose. We define Pose Parameter Prror (PPE) as the $L_{2}$ distance between the generated and ground truth 6-dimensional head pose parameters, which measures the accuracy of the head movements. To measure the alignment between the input singing audio and head movements, we adopt the Beat Align Score (BA) [88] to the generated head pose parameters. (3) Diversity. To measure the diversity of the generation results, we calculate the average $L_{2}$ distance of 68 facial landmarks between all pairs from the generated $n$ samples, which can be written as $\frac{1}{n(n-1)}{\textstyle\sum_{i=1}^{n}}{\textstyle\sum_{j\neq i}^{n}}||\boldsymbol{L}_{i}-\boldsymbol{L}_{j}||_{2}^{2}$. (4) Inference speed. It is important to generate 3D facial motion in real time according to the input audio. Therefore, we measure the model inference speed by calculating the average number of frames generated per second, termed FPS.
Baseline methods. We select some typical 3D facial animation methods as the baseline methods, including FaceFormer [4], CodeTalker [5], Imitator [12], SelfTalk [13] and FaceDiffuser [14]. We not only evaluate their pre-trained models on the SingingHead test set, but also train them on our SingingHead dataset. Since most of these methods [4, 5, 12, 13] are designed for mesh sequences, while the 3D facial motion in our dataset is represented by FLAME parameters [35], so we use the FLAME mesh sequences without head pose to train these methods. FaceDiffuser [14] and the 3D facial animation module of our UniSinger framework are trained directly using the FLAME parameters which include the head pose parameters.
The quantitative and qualitative results are shown in Tab. II and Fig. 11, respectively. On the one hand, pre-trained models show the worst quantitative and qualitative results, which indicates that the 3D facial animation models trained on talking datasets cannot generalize well to singing audio. After training on the SingingHead dataset, almost all the metrics and visualization results of these models have improved, which demonstrates the necessity of our singing-specific dataset for the 3D singing head animation task. On the other hand, looking at all the models that are trained on the SingingHead dataset, our method achieves more realistic and vivid visual results, and competitive results on all metrics. Specifically, nearly all the previous methods [4, 5, 12, 13] use offsets of mesh vertices to represent the facial movements, which is not only computationally intensive but also hard to generate head movements. In contrast, our method uses sequences of the 59-dimensional FLAME parameters [35] including head pose parameters to represent motions, leading to more accurate head movements and vivid results as evidenced by the PPE and BA metrics. Moreover, most of the existing methods [4, 5, 12, 13] model the audio-driven 3D facial animation as a deterministic regression task, which can only generate one results given input audio. Such methods can regress more accurate vertex positions, leading to better vertex-level metrics such as LVE and FVE. In contrast, FaceDiffuser [14] and our method model it as a generation task, which can generate diverse facial motions according to one input audio. Generative models tend to produce inferior results on deterministic metrics. Nevertheless, our method achieves competitive results on LVE and FVE metrics and demonstrates some diversity in generation as shown in Fig. 9. Achieving accuracy and diversity in generation involves a trade-off. Although our method exhibits less diversity compared to FaceDiffuser [14], it achieves much better lip and face synchronization. In addition, our method has the fastest inference speed and can achieve real-time 3D singing head animation, thanks to the non-autoregressive mechanism we employed.
We also conduct ablation studies on the proposed 3D facial animation module of the UniSinger framework as shown in Tab. III. The injection of background music helps the model generate more accurate facial and head movements, while slightly reducing the variety of generated results due to the additional conditions introduced. Reconstruction loss $\mathcal{L}_{rc}$ and KL loss $\mathcal{L}_{kl}$ are necessary for the training of VAE. Without $\mathcal{L}_{rc}$, the latent space of the model tends to concentrate in a small region of the prior distribution, resulting in limited diverse generation ability. Without $\mathcal{L}_{kl}$, the model is hard to converge, leading to poor accuracy and diversity in generation. Velocity loss $\mathcal{L}_{vel}$ is used to prevent the generation of static or motionless results. Mesh vertices loss $\mathcal{L}_{mesh}$ contributes to increasing the realism and diversity of the generated results, as well as improving the alignment with the input audio.
TABLE III: Ablation study for vae-based 3D facial animation module. The units of LVE, FVE, and FDD are $mm$, $mm$, and ×$\times 10^{-2}mm$, respectively.
Method LVE $\downarrow$ FVE $\downarrow$ FDD $\downarrow$ BA $\uparrow$ PPE $\downarrow$ Diversity $\uparrow$
w/o BGM 4.8025 1.1017 1.1874 0.2343 0.1042 1.2407
w/o $L_{rc}$ 4.4723 1.0035 1.4530 0.2288 0.1475 0.0042
w/o $L_{kl}$ 4.9272 1.1274 0.8862 0.2402 0.1031 0.5593
w/o $L_{vel}$ 4.5952 1.0397 1.1991 0.2389 0.0966 0.9608
w/o $L_{mesh}$ 4.5611 1.0230 1.2532 0.2355 0.0933 0.7873
Ours 4.5928 1.0435 1.2286 0.2378 0.0988 1.0993
Figure 9: Illustration of the diverse generation. Our model can generate diverse 3D singing head motions and corresponding portrait videos from the same input singing audio.
TABLE IV: Quantitative results of 2D singing portrait video synthesis. We report the evaluation metrics on the generated singing videos of the selected 10 identities. We list results under the resolution of 256 / 512, respectively.
Method FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CPBD $\uparrow$ M-LMD $\downarrow$ F-LMD $\downarrow$ BA $\uparrow$
Ground Truth 0.000 0.000 1.000 0.401 / 0.299 0.000 0.000 0.252 / 0.249
MakeItTalk [25] 29.21 / 46.90 0.180 / 0.276 0.620 / 0.690 0.226 / 0.034 2.753 / 5.419 4.510 / 8.972 0.243 / 0.238
Audio2Head [15] 42.06 / 62.98 0.233 / 0.328 0.560 / 0.647 0.228 / 0.044 2.711 / 5.343 5.002 / 10.01 0.249 / 0.254
SadTalker [1] 18.07 / 18.73 0.175 / 0.238 0.592 / 0.660 0.217 / 0.161 2.989 / 5.887 4.774 / 9.508 0.237 / 0.239
EAT [16] 54.28 / 75.58 0.238 / 0.340 0.553 / 0.640 0.230 / 0.048 2.672 / 5.248 5.112 / 10.18 0.244 / 0.243
DreamTalk [17] 53.91 / 66.25 0.297 / 0.404 0.492 / 0.592 0.245 / 0.048 3.655 / 7.221 8.261 / 16.42 0.221 / 0.219
UniSinger (2D) 12.31 / 12.08 0.169 / 0.233 0.618 / 0.688 0.258 / 0.165 2.041 / 3.991 3.527 / 7.040 0.251 / 0.251